For numbers 2-4

#2 When abdominoplasty costs were $6,000, Dr. Hurwitz treated 300 patients.

When they increased the price to $7,400,customers decreased to 250. What is the elasticity of demand for abdominoplasty using arc elasticity or the midpoints formula (rounded-off to 2 decimal places)?

A. - 0.21

B. - 0.87

C. - 2.03

D. -1.15

E. -8.70

#3 The elasticity of demand for abdominoplasty can be described as

A. elastic

B. inelastic

C. unitary elastic

D.all of the above

#4. In terms of raising revenue (or sales) as a goal

A.The doctor was correct in increasing price

B. The doctor was incorrect in raising price

C.Revenue did not change with the doctor increasing price

D.none of the above
Answer #1

a. Mid-point elasticity of demand is given as below:

ED =

ED =

ED = -50/275 1400/6700

ED = -0.1818 0.2089

ED = -0.87

Option B is correct.

b. Option B is correct.

The elasticity value of less than 1 indicates that the demand is inelastic. With the change in price, the proportionate change in demand is less.

c. Option A is correct.

The higher would be the revenue if the demand is inelastic and the price is increased.

Initial revenue = 6000 * 300 = $90000

Final revenue = 7400 * 250 = $1850000.